### v1

# Ondřej Mokrý Dept. of Telecommunications

Abstract
The paper focuses on inpainting missing parts of an audio signal spectrogram. First, a recent successful approach based on an untrained neural network is revised and its several modifications are proposed, improving the signal-to-noise ratio of the restored audio. Second, the Janssen algorithm, the autoregression-based state-of-the-art for time-domain audio inpainting, is adapted for the time-frequency setting. This novel method, coined Janssen-TF, is compared to the neural network approach using both objective metrics and a subjective listening test, proving Janssen-TF to be superior in all the considered measures.
Index Terms:
audio inpainting, autoregression, deep prior, spectrogram.
I Introduction
Audio inpainting, sometimes referred to as missing audio interpolation or concealment, is the task of filling in the absent parts of audio with a meaningful content. Typically, the missing information is in the form of lost samples in the time domain. If compact regions of such samples are present, they are often referred to as the gaps. For the described task, plenty of methods have been developed: those based on the autoregressive character of audio [1, 2], methods based on the sparsity of the time-frequency representation of audio [3, 4, 5, 6], or methods relying on similarities in audio recordings [7, 8, 9]. Quite recently, deep learning audio inpainting methods appeared [10, 11]. For gaps of up to 50 ms in length, the Janssen method [2, 12] has been evaluated the best for a long time, but PHAIN [13] may become a new state-of-the art.
The time-domain inpainting represents the most natural case from the point of view of practical applications (signal recovery). Nonetheless, many practical algorithms, such as the classic audio coding algorithms (MP3, AAC, etc.), operate on short-time signal spectra. From this perspective, it is valuable to think about inpainting in the time-frequency domain. [^1] In other fields, one can go even further and require inpainting purely in the frequency domain, see for example [14], [15]. In such a case, a portion of time-frequency (TF) coefficients222 typically in the form of a complex-valued matrix are missing, which need to be recovered to allow a later reconstruction of the time-domain signal. Indeed, there have been a few methods developed in the TF context, all of them relying on neural modeling [16, 17, 18]. The most recent work [18] is based on the concept of a “deep prior”, a neural network that does not require training in the common sense.
Note that a compact gap in time bears some fundamental differences compared with a gap in the form of missing a number of consecutive columns of a spectrogram. This follows from the nature of time-frequency transforms, which use windows with overlaps to compute spectrograms, see for instance [19]. In particular, a gap in a spectrogram does not fully correspond to a time-domain gap and vice versa; there is always an amplitude fading in one of the domains.
The goal of this paper is to adopt the Janssen algorithm [2], proven successful in the time-domain case, to the TF-domain case, and compare its performance with the recent deep-prior-based approach of [18], both using objective metrics and by listening tests.
II Deep Prior Audio Inpainting (DPAI)
The deep prior, introduced in [20], is a concept used for the reconstruction of corrupted signals using an “untrained” convolutional neural network (CNN). The signal reconstruction involves only a proper CNN architecture and the corrupted signal. The deep prior concept has also proven that the network architecture matters much more than previously thought.
II-A Application of Deep Prior to Audio Inpainting
Deep prior reconstruction can be utilized in a variety of applications [20]. In the context of the time-domain audio inpainting, consider a corrupted signal $\mathbf{x}^{\textup{cor}}\!\in\mathbb{R}^{N}$, having missing regions (filled with zeros), compared with the original signal $\mathbf{x}^{\textup{orig}}\!\in\mathbb{R}^{N}$. Formally, if a binary mask $\mathbf{m}\in\{0,1\}^{N}$ is utilized to represent the missing sections of the signal, the relation $\mathbf{x}^{\textup{cor}}=\mathbf{m}\odot\mathbf{x}^{\textup{orig}}$ holds, where $\odot$ represents the elementwise product. Such a relation naturally characterizes the feasible set $\mathit{\Gamma}_{\textup{T}}=\{\mathbf{x}\,|\,\mathbf{m}\odot\mathbf{x}=\mathbf{m}\odot\mathbf{x}^{\textup{orig}}\}$.
An encoder–decoder CNN denoted $f_{\theta}(\mathbf{n})$ is assumed to be able to generate any signal of size $N$ using a random but fixed noise input $\mathbf{n}\in\mathbb{R}^{N}$, using a proper $\theta$. This is a reliable assumption since the number of network parameters ${\theta}$ (weights and biases) is significantly greater than $N$.
In the deep prior reconstruction, parameters $\theta$ are initialized randomly. To obtain a reconstruction result, $\theta$ is then optimized as follows:
$\theta^{*}=\operatorname*{argmin}_{\theta}E(\mathbf{m}\odot f_{\theta}(\mathbf{n}),\mathbf{x}^{\textup{cor}}),$ (1)
where $E$ is a differentiable data fidelity function measuring the closeness of its two inputs. The optimal $\theta^{*}$ can be obtained using an iterative optimizer, e.g., Adam [21]. The crucial observation is that the optimization must be interrupted before it reaches $\theta^{*}$, for which we would have $f_{\theta^{*}\!}(\mathbf{n})=\mathbf{x}^{\textup{cor}}$. Instead, early stopping of the training process is desired, resulting in an estimate $\hat{\theta}$ and a corresponding $\hat{\mathbf{x}}=f_{\hat{\theta}}(\mathbf{n})$ [20].
The authors of [18] used an analog approach to inpaint audio in the transformed domain. Through the use of the short-time Fourier transform (STFT), an audio recording $\mathbf{x}$ is considered merely in the TF domain, where it forms a spectrogram matrix $\mathbf{X}\in\mathbb{C}^{T\times F}$. Therefore the CNN operates on spectrograms and $\hat{\mathbf{X}}=f_{\hat{\theta}}(\mathbf{N})$ for a random but fixed matrix $\mathbf{N}$. Once $\hat{\mathbf{X}}$ has been established, the time-domain solution is obtained by the inverse STFT.
Regarding the mask, in [18] only the case of completely missing columns of the spectrogram is considered, i.e., for the TF mask $\mathbf{M}\in\{0,1\}^{T\times F}$ it holds that each column is either full of ones or zeros. Nevertheless, generalization to other TF masks is straightforward. The observed spectrogram $\mathbf{X}^{\textup{cor}}$ is finally modeled as $\mathbf{X}^{\textup{cor}}\!=\mathbf{M}\odot\mathbf{X}^{\textup{orig}}$, where $\mathbf{X}^{\textup{orig}}$ corresponds to $\mathbf{x}^{\textup{orig}}$ in the time domain. Note that by analogy with the time-domain case, we have the feasible set $\mathit{\Gamma}_{\textup{TF}}=\{\mathbf{X}\,|\,\mathbf{M}\odot\mathbf{X}=\mathbf{M}\odot\mathbf{X}^{\textup{orig}}\}$.
The loss function $E$ in (1), which is optimized with respect to $\theta$, is in the form
$E=\alpha_{1}\mathrm{MSE}(\mathbf{M}\odot\mathbf{X},\mathbf{X}^{\textup{cor}})+\alpha_{2}\mathrm{MSS}(\mathbf{M}\odot\mathbf{X},\mathbf{X}^{\textup{cor}}),$ (2)
where $\mathrm{MSS}$ is the multi-scale spectrogram loss [22]. In [18], the weights $\alpha_{1}$ and $\alpha_{2}$ have values 1 and 0.1 respectively.
II-B Architecture
The U-Net is a well-known CNN architecture consisting of several convolutional layers, downsampling (encoder) and upsampling (decoder) operations. In certain applications including audio inpainting, the existence of skip connections between the encoder and decoder stages is important. The authors of [18] used an architecture similar to U-Net called MultiResUNet [23] with an additional inclusion of the so-called harmonic convolution [24], which proved to be the best among multiple tested architectures in [18]. The harmonic convolution is a variation on the common 2D convolution that takes into account higher harmonics of audio.
II-C Modifications
The DPAI reconstructs the whole $\hat{\mathbf{X}}$. Yet, considering that only a part of the spectrogram is not known, the authors of [18] evaluate the performance of their method after substituting the coefficients of $\hat{\mathbf{X}}$ that correspond to ones in the mask $\mathbf{M}$ with the observed data. Formally, such an action is a projection
$\operatorname{proj}_{\mathit{\Gamma}_{\textup{TF}}}(\hat{\mathbf{X}})=\mathbf{M}\odot\mathbf{X}^{\textup{cor}}+(\mathbf{1}-\mathbf{M})\odot\hat{\mathbf{X}}$ (3)
onto the set $\mathit{\Gamma}_{\textup{TF}}$. Note that a resulting sharp transition between the regions is smeared in the time domain due to the nature of the STFT [19]. We call the modified reconstructions with context, while the others using the entire $\hat{\mathbf{X}}$ are without context.
The other modifications come from the fact that the reconstruction always starts from a random initialization of both $\theta$ and $\mathbf{N}$. As a consequence, the objective and the subjective qualities of the reconstructed signal might slightly vary, based on the particular realization. We observed that averaging the reconstructions from multiple independent reconstructions of the same signal improves the objective performance metrics. Obviously, the cost of such an approach grows linearly with the number of trials. We also noticed that a slight improvement can be assured by averaging a number of final epochs of an individual reconstruction.
In the experimental part, we will study whether the proposed modification of DPAI brings a significant performance improvement.
III Janssen spectrogram inpainting (Janssen-TF)
A signal $\mathbf{x}=[x_{1},\dots,x_{N}]^{\top}\!\in\mathbb{R}^{N}$ corresponds to an autoregressive process if the following holds:
$\sum_{i=1}^{p+1}a_{i}x_{n+1-i}=e_{n},\quad n=1,\dots,N+p,$ (4)
where $\mathbf{a}=[1,a_{2},\dots,a_{p+1}]^{\top}$ is the vector of model parameters (AR coefficients), $p$ is the model order, and $\mathbf{x}$ is zero-padded such that its final length is $N+p$. The error term $\mathbf{e}\in\mathbb{R}^{N+p}$ represents a realization of a zero-mean white noise process with variance $\sigma^{2}$ [25, Def. 3.1.2].
In the inpainting method proposed by Janssen et al. [2], the restored audio signal is supposed to follow the AR model (4) while both the missing samples in the time-domain and the unknown AR parameters $\mathbf{a}$ are being estimated. Janssen et al. approach the problem iteratively by repeating two principal steps in the $i$-th iteration:
  1. 1.
    for a temporary solution $\mathbf{x}^{(i)}$, estimate the AR model parameters $\mathbf{a}^{(i)}$;
  2. 2.
    given the model parameters $\mathbf{a}^{(i)}$, estimate a new temporary solution $\mathbf{x}^{(i+1)}$ which fits the observed samples and the norm of the model error $\mathbf{e}$ is minimized.
Step 1 can be solved efficiently using standard methods, such as the Durbin–Levinson recursion (see e.g. the function lpc in Matlab). Regarding step 2, we propose to adapt the Janssen iteration to the TF setting, denoted Janssen-TF. In our adaptation, the signal spectrogram is constrained to fit the observation, $\mathbf{X}^{\textup{cor}}$, while the signal follows the estimated AR model. To formalize this modification, note that the error in (4) can be written as $\mathbf{e}=\mathbf{A}\mathbf{x}$, where $\mathbf{A}\in\mathbb{R}^{(N+p)\times N}$ is a Toeplitz matrix composed using the (fixed) coefficients $\mathbf{a}$. We can then pose the proposed modification of step 2 as
$\mathbf{x}^{(i+1)}=\arg\min_{\mathbf{x}}\tfrac{1}{2}\|\mathbf{A}\mathbf{x}\|^{2}+\iota_{\mathit{\Gamma}_{\textup{TF}}}(\mathbf{L}\mathbf{x}),$ (5)
where $\iota_{\mathit{\Gamma}_{\textup{TF}}}$ denotes the indicator function of the convex set $\mathit{\Gamma}_{\textup{TF}}$ of feasible spectrograms. [^3] Note that the original Janssen algorithm minimizes the function $\frac{1}{2}\|\mathbf{A}\mathbf{x}\|+\iota_{\mathit{\Gamma}_{\textup{T}}}(\mathbf{x})$ in subproblem 2. As such, the problem (5) is convex and it is solvable using standard numerical methods. We propose to employ the alternating direction method of multipliers (ADMM), which includes two principal steps [26, Sec. 3.1.1]:
$\bar{\mathbf{x}}^{(k+1)}$ $=\arg\min_{\mathbf{x}}\tfrac{1}{2}\|\mathbf{A}\mathbf{x}\|^{2}+\tfrac{\rho}{2}\|\mathbf{L}\mathbf{x}-\mathbf{z}^{(k)}+\mathbf{u}^{(k)}\|^{2}\!,$ (6a)
$\mathbf{z}^{(k+1)}$ $=\arg\min_{\mathbf{z}}\iota_{\mathit{\Gamma}_{\textup{TF}}}(\mathbf{z})+\tfrac{\rho}{2}\|\mathbf{L}\bar{\mathbf{x}}^{(k+1)}-\mathbf{z}+\mathbf{u}^{(k)}\|^{2}\!.$ (6b)
The whole algorithm is summarized in Alg. 1. For the simplification of (6a), we have posed the assumption [^4] This assumption is met for instance if $\mathbf{L}$ corresponds to a Parseval tight frame, which can be ensured by a suitable setting of the STFT. $\mathbf{L}^{*}\mathbf{L}=\mathbf{I}$, leading to $\bar{\mathbf{x}}^{(k+1)}=\operatorname{prox}_{\frac{1}{2}\|\cdot\|^{2}}(\mathbf{L}^{*}(\mathbf{z}^{(k)}-\mathbf{u}^{(k)}))$, see [27, Rem. 3], which then corresponds to line 1 of Alg. 1 [28, Table 1]. Similarly, (6b) boils down to line 1 of the algorithm; this is computed via (3). The hyperparameters of the algorithm are the number of outer and inner iterations ($I$ and $K$, respectively), and the step size $\rho>0$.
1 initialize $\mathbf{x}^{(1)}$
2 for $i=1,\dots,I$ do
       // AR model update
3       $\mathbf{a}^{(i)}=\texttt{lpc}(\mathbf{x}^{(i)})$
       // signal update using ADMM
4       form the Toeplitz matrix $\mathbf{A}$ from $\mathbf{a}^{(i)}$
5       $\mathbf{z}^{(1)}=\mathbf{L}\mathbf{x}^{(i)}$, $\mathbf{u}^{(1)}=\mathbf{0}$
6       for $k=1,\dots,K$ do
7             $\bar{\mathbf{x}}^{(k+1)}=(\mathbf{I}+\frac{1}{\rho}\mathbf{A}\!^{\top}\!\mathbf{A})^{-1}\mathbf{L}^{*}(\mathbf{z}^{(k)}-\mathbf{u}^{(k)})$
8             $\mathbf{z}^{(k+1)}=\operatorname{proj}_{\mathit{\Gamma}_{\textup{TF}}}(\mathbf{L}\bar{\mathbf{x}}^{(k+1)}+\mathbf{u}^{(k)})$
9             $\mathbf{u}^{(k+1)}=\mathbf{u}^{(k)}+\mathbf{L}\bar{\mathbf{x}}^{(k+1)}-\mathbf{z}^{(k+1)}$
10       end for
11      $\mathbf{x}^{(i+1)}=\bar{\mathbf{x}}^{(K+1)}$
12
13 end for
Algorithm 1 Janssen-TF-ADMM
Note that several algorithms can be used to solve (5) numerically, such as the primal–dual algorithm [29] or the Douglas–Rachford or the forward–backward algorithms [28], combined with the approximal operator of the function $\iota_{\mathit{\Gamma}_{\textup{TF}}}(\mathbf{L}\,\cdot)$ [30]. However, the variant with ADMM described above, denoted Janssen-TF-ADMM, exhibited the best performance out of six algorithms in our preliminary testing.
IV Experiments
IV-A Setup and metrics
For our experiments, we adopt the audio dataset [^5] https://github.com/fmiotello/dpai/ used for the evaluation of DPAI [18]. It consists of 6 music and 2 speech recordings of 5 seconds in length, sampled at 16 kHz. Only minor modifications were made to the original distribution of gaps (i.e., the placement of the missing spectrogram columns), such that each test excerpt includes 5 gaps of the same length. The gap lengths range from 1 up to 6 missing columns. In turn, the entire test set comprises of $8\times 6=48$ corrupted audio signals. Even though the dataset is relatively small and includes only a few types of audio signals, we decided to stick to it because its use justifies a fair comparison of the proposed method with DPAI. The choice allows the use of exactly the same U-net architecture of DPAI from [18]. In particular, we choose their “best2” setting. The number of epochs of the Adam optimizer is 5000.
The STFT (i.e., the operator $\mathbf{L}$) utilizes the 2048-sample-long Hann window (which corresponds to 128 ms), with 75% overlap and 2048 frequency channels; this setting is the default for the time-frequency transform of the librosa package [^6] https://librosa.org/doc/0.10.2/generated/librosa.stft.html, as used in [18], and ensures that $\mathbf{L}$ forms a Parseval tight frame.
To objectively evaluate the reconstruction quality, we assess the difference between the inpainted signal $\hat{\mathbf{x}}$ and a clean reference $\mathbf{x}^{\textup{orig}}$ using two standard metrics:
  1. 1.
    signal-to-noise ratio (SNR), defined in decibels as $\text{SNR}(\hat{\mathbf{x}},\mathbf{x}^{\textup{orig}})=10\log_{10}\frac{\|\mathbf{x}^{\textup{orig}}\|^{2}}{\|\mathbf{x}^{\textup{orig}}-\hat{\mathbf{x}}\|^{2}}$ [3],
  2. 2.
    objective difference grade (ODG) computed using the PEMO-Q [31], which takes psychoacoustics into account and predicts the subjective evaluation of the difference between $\hat{\mathbf{x}}$ and $\mathbf{x}^{\textup{orig}}$ on the scale from $-4$ to 00 (from “very annoying” to “imperceptible”).
Above that, we run
  1. 3.
    a subjective listening test containing a limited number of excerpts. The MUSHRA-type test [32] was performed that includes the reference signal, the DPAI reconstruction with context, the DPAI reconstruction without context, the DPAI with context but averaged result from four random initializations, the gapwise Janssen method in the time domain [12], the Janssen-TF-ADMM, and finally the hidden anchor, i.e., the signal reconstructed from the spectrogram with gaps.
The participants were asked to score the perceived similarity of the signals with the reference on a continuous scale from 0 to 100, with the help of the webMUSHRA environment [^7] https://github.com/audiolabs/webMUSHRA/. The test was run in a quiet music studio, using a professional sound card and headphones. The listening conditions were identical for all the participants. Following the recommendation [32], 11 assessors out of 18 were selected through post-screening.
IV-B Results
First, we evaluate the proposed modifications of DPAI. As seen in Fig. 1, averaging the realizations brings improvement in quality in terms of SNR, but it comes at the cost of a higher time complexity. The differences observed in Fig. 1 were further tested using the Wilcoxon signed rank test [^8] https://www.mathworks.com/help/stats/signrank.html, which proved that the median of the SNR using DPAI with averaging is significantly higher compared to the original DPAI with context. On the other hand, a significant difference in case of the ODG metric has not been proven.
001010101020202020303030304040404050505050001010101020202020303030304040404050505050SNR (dB), DPAI with contextSNR (dB), DPAI averagedGaplength1 2 3 4 5 6 44-4- 433-3- 322-2- 211-1- 10044-4- 433-3- 322-2- 211-1- 100ODG, DPAI with contextODG, DPAI averaged
Figure 1: Comparison of DPAI reconstruction with context versus the proposed DPAI variant averaging four random initializations.
In contrast to the averaging of realizations, the averaging across the final epochs from a single realization does not increase the computational load and brings better quality mostly to audio recordings reconstructed without context—the greatest improvement (averaged over test signals) was 0.16 ODG for the gap of 2 columns and 15 epochs averaged. However, this strategy has a negligible impact when considering reconstructions with context.
The objective comparison that includes the proposed Janssen-TF method is shown in Fig. 2. The first observation is that the time-domain-based Janssen method [12] cannot compete with the other methods. This is due to the incompatibility of gaps in the time and time-frequency domains (a sharp mask in the spectrogram domain affects a longer signal portion when transformed into the time domain). While the average performance of the DPAI variants mostly lie within the confidence intervals of each other, suggesting limited significance of the differences, the proposed Janssen-TF-ADMM method largely outperforms the competitors in both SNR and ODG (except for the longest gaps).
Figure 2: Comparison of the inpainting methods using objective metrics – SNR (top) and ODG (bottom). The plots show results averaged over the 8 test signals together with the bootstrap interval estimates of the mean values at the $\alpha=5\%$ significance level [33, p. 160]. If the intervals do not overlap, it may be concluded that the difference of the means is statistically significant.
To support the objective evaluation, Fig. 3 presents the results of the subjective listening test. Its results correlate well with the objective evaluation; in particular, Janssen-TF-ADMM is superior to the competing methods. Moreover, this result is statistically significant in terms of medians. On the other hand, the DPAI variants scored almost the same. This is in line with the ODG results in Fig. 1, even though the listening test featured only a selection of the test signals.
anchorreferenceDPAI without contextDPAI with contextDPAI averagedgapwise JanssenJanssen-TF-ADMM0020202020404040406060606080808080100100100100MUSHRA score
Figure 3: A boxplot showing the distribution of scores in the listening test. The individual boxes span from the 25th to the 75th percentile of the recorded scores. The notches (filled areas) around the medians (orange lines) are constructed such that boxes whose notches do not overlap have different medians at the 5% statistical significance level.
A single run of DPAI takes ca 19 minutes on the NVIDIA Tesla V100S GPU with 32 GB of memory, regardless the number of missing spectrogram columns. Computing a single signal with Janssen-TF-ADMM takes 10 to 20 minutes, directly proportional to the number of columns; this corresponds to a PC with Intel Core i7 3.40 GHz processor, 32 GB RAM.
V Conclusion and Outlook
The paper has shown that the proposed method of spectrogram inpainting, Janssen-TF, performs significantly better than the recently introduced DPAI algorithm, which is based on the deep prior idea. This conclusion has been certified both by objective and subjective tests. The proposed minor modifications of DPAI led to a slight improvement in terms of the SNR, but have not been reflected in the perceived sound quality.
In the future, it might be interesting to use the same adaptation strategy with other methods for the time-domain inpainting, such as the PHAIN regularized algorithm [13].
Matlab source codes for Janssen-TF and other supplementary material are publicly available at https://github.com/rajmic/spectrogram-inpainting.

### v2

@@
 The paper focuses on inpainting missing parts of an audio signal spectrogram.
-First, a recent successful approach based on an untrained neural network is revised and its several modifications are proposed, improving the signal-to-noise ratio of the restored audio.
-Second, the Janssen algorithm, the autoregression-based state-of-the-art for time-domain audio inpainting, is adapted for the time-frequency setting.
-This novel method, coined Janssen-TF, is compared to the neural network approach using both objective metrics and a subjective listening test, proving Janssen-TF to be superior in all the considered measures.
+The autoregression-based Janssen algorithm, the state-of-the-art for the time-domain audio inpainting, is adapted for the time-frequency setting.
+This novel method, termed Janssen-TF, is compared to the deep-prior neural network approach using both objective metrics and a subjective listening test, proving Janssen-TF to be superior in all the considered measures.
 Index Terms:
-audio inpainting, autoregression, deep prior, spectrogram.
+audio inpainting, autoregression, deep prior, time-frequency, spectrogram.
 I Introduction
@@
 For gaps of up to 50 ms in length, the Janssen method [2, 12] has been evaluated the best for a long time, but PHAIN [13] may become a new state-of-the art.
-The time-domain inpainting represents the most natural case from the point of view of practical applications (signal recovery).
-Nonetheless, many practical algorithms, such as the classic audio coding algorithms (MP3, AAC, etc.), operate on short-time signal spectra.
+The time-domain inpainting represents the most natural case from the point of view of the application to signal recovery.
+Nonetheless, many practical algorithms, such as classic audio codecs (MP3, AAC, etc.), operate on short-time signal spectra.
 From this perspective, it is valuable to think about inpainting in the time-frequency domain.
@@
 Indeed, there have been a few methods developed in the TF context, all of them relying on neural modeling [16, 17, 18].
-The most recent work [18] is based on the concept of a “deep prior”, a neural network that does not require training in the common sense.
-Note that a compact gap in time bears some fundamental differences compared with a gap in the form of missing a number of consecutive columns of a spectrogram.
+The most recent work [18] is based on the concept of a “deep prior”, a neural network that does not need training in the most common sense.
+Note that a compact gap in time bears some fundamental differences compared with a gap in the form of a missing number of consecutive columns of a spectrogram.
 This follows from the nature of time-frequency transforms, which use windows with overlaps to compute spectrograms, see for instance [19].
-In particular, a gap in a spectrogram does not fully correspond to a time-domain gap and vice versa; there is always an amplitude fading in one of the domains.
-The goal of this paper is to adopt the Janssen algorithm [2], proven successful in the time-domain case, to the TF-domain case, and compare its performance with the recent deep-prior-based approach of [18], both using objective metrics and by listening tests.
+In particular, a gap in a spectrogram does not fully correspond to a time-domain gap and vice versa; there is always an amplitude fading in one of the domains thanks to the windowing.
+The goal of this paper is to adapt the Janssen algorithm [2], proven successful in the time-domain case, for the TF-domain case, and compare its performance with the recent deep-prior-based approach of [18], both using objective metrics and by listening tests.
 II Deep Prior Audio Inpainting (DPAI)
-The deep prior, introduced in [20], is a concept used for the reconstruction of corrupted signals using an “untrained” convolutional neural network (CNN).
-The signal reconstruction involves only a proper CNN architecture and the corrupted signal.
+The deep prior, introduced in [20], is a concept used for the reconstruction of corrupted signals using a convolutional neural network (CNN).
+In contrast to most CNN applications, [20] does not utilize a training dataset; the reconstruction only involves a proper CNN architecture and the corrupted signal.
 The deep prior concept has also proven that the network architecture matters much more than previously thought.
 II-A Application of Deep Prior to Audio Inpainting
-Deep prior reconstruction can be utilized in a variety of applications [20].
+The deep prior reconstruction can be utilized in a variety of applications, as outlined in [20].
 In the context of the time-domain audio inpainting, consider a corrupted signal $\mathbf{x}^{\textup{cor}}\!\in\mathbb{R}^{N}$, having missing regions (filled with zeros), compared with the original signal $\mathbf{x}^{\textup{orig}}\!\in\mathbb{R}^{N}$.
@@
 The optimal $\theta^{*}$ can be obtained using an iterative optimizer, e.g., Adam [21].
-The crucial observation is that the optimization must be interrupted before it reaches $\theta^{*}$, for which we would have $f_{\theta^{*}\!}(\mathbf{n})=\mathbf{x}^{\textup{cor}}$.
-Instead, early stopping of the training process is desired, resulting in an estimate $\hat{\theta}$ and a corresponding $\hat{\mathbf{x}}=f_{\hat{\theta}}(\mathbf{n})$ [20].
+The crucial observation is that the optimization must be interrupted before it reaches $\theta^{*}$, for which we would have $f_{\theta^{*}\!}(\mathbf{n})=\mathbf{x}^{\textup{cor}}$, meaning basically overfitting.
+Instead, early stopping of the training is desired, resulting in an estimate $\hat{\theta}$ and a corresponding $\hat{\mathbf{x}}=f_{\hat{\theta}}(\mathbf{n})$ [20].
 The authors of [18] used an analog approach to inpaint audio in the transformed domain.
-Through the use of the short-time Fourier transform (STFT), an audio recording $\mathbf{x}$ is considered merely in the TF domain, where it forms a spectrogram matrix $\mathbf{X}\in\mathbb{C}^{T\times F}$.
+Through the use of the short-time Fourier transform (STFT), denoted $\mathbf{L}$ in the following, an audio recording $\mathbf{x}$ is considered merely in the TF domain, where it forms a spectrogram matrix $\mathbf{L}\mathbf{x}=\mathbf{X}\in\mathbb{C}^{T\times F}$.
 Therefore the CNN operates on spectrograms and $\hat{\mathbf{X}}=f_{\hat{\theta}}(\mathbf{N})$ for a random but fixed matrix $\mathbf{N}$.
@@
 Regarding the mask, in [18] only the case of completely missing columns of the spectrogram is considered, i.e., for the TF mask $\mathbf{M}\in\{0,1\}^{T\times F}$ it holds that each column is either full of ones or zeros.
-Nevertheless, generalization to other TF masks is straightforward.
-The observed spectrogram $\mathbf{X}^{\textup{cor}}$ is finally modeled as $\mathbf{X}^{\textup{cor}}\!=\mathbf{M}\odot\mathbf{X}^{\textup{orig}}$, where $\mathbf{X}^{\textup{orig}}$ corresponds to $\mathbf{x}^{\textup{orig}}$ in the time domain.
+Nevertheless, a generalization to other TF masks is straightforward.
+The observed spectrogram $\mathbf{X}^{\textup{cor}}$ is finally modeled as $\mathbf{X}^{\textup{cor}}\!=\mathbf{M}\odot\mathbf{X}^{\textup{orig}}$, where $\mathbf{X}^{\textup{orig}}$ corresponds to $\mathbf{x}^{\textup{orig}}$ in the time domain, i.e., $\mathbf{X}^{\textup{orig}}=\mathbf{L}\mathbf{x}^{\textup{orig}}$.
 Note that by analogy with the time-domain case, we have the feasible set $\mathit{\Gamma}_{\textup{TF}}=\{\mathbf{X}\,|\,\mathbf{M}\odot\mathbf{X}=\mathbf{M}\odot\mathbf{X}^{\textup{orig}}\}$.
@@
 $E=\alpha_{1}\mathrm{MSE}(\mathbf{M}\odot\mathbf{X},\mathbf{X}^{\textup{cor}})+\alpha_{2}\mathrm{MSS}(\mathbf{M}\odot\mathbf{X},\mathbf{X}^{\textup{cor}}),$ (2)
-where $\mathrm{MSS}$ is the multi-scale spectrogram loss [22]. In [18], the weights $\alpha_{1}$ and $\alpha_{2}$ have values 1 and 0.1 respectively.
+where MSE is the Mean Squared Error and $\mathrm{MSS}$ is the multi-scale spectrogram loss [22]. In [18], the weights $\alpha_{1}$ and $\alpha_{2}$ have values 1 and 0.1 respectively.
 II-B Architecture
@@
 The harmonic convolution is a variation on the common 2D convolution that takes into account higher harmonics of audio.
-II-C Modifications
+II-C Variants
 The DPAI reconstructs the whole $\hat{\mathbf{X}}$.
-Yet, considering that only a part of the spectrogram is not known, the authors of [18] evaluate the performance of their method after substituting the coefficients of $\hat{\mathbf{X}}$ that correspond to ones in the mask $\mathbf{M}$ with the observed data.
+Yet, considering that only a part of the spectrogram was unavailable, the authors of [18] evaluate the performance of their method after injecting the reliable coefficients back to the solution as a postprocessing step.
 Formally, such an action is a projection
@@
 Note that a resulting sharp transition between the regions is smeared in the time domain due to the nature of the STFT [19].
-We call the modified reconstructions with context, while the others using the entire $\hat{\mathbf{X}}$ are without context.
-The other modifications come from the fact that the reconstruction always starts from a random initialization of both $\theta$ and $\mathbf{N}$.
-As a consequence, the objective and the subjective qualities of the reconstructed signal might slightly vary, based on the particular realization.
-We observed that averaging the reconstructions from multiple independent reconstructions of the same signal improves the objective performance metrics.
-Obviously, the cost of such an approach grows linearly with the number of trials.
-We also noticed that a slight improvement can be assured by averaging a number of final epochs of an individual reconstruction.
-In the experimental part, we will study whether the proposed modification of DPAI brings a significant performance improvement.
+We call the reconstructions modified by (3) with context, while the others using the entire $\hat{\mathbf{X}}$ are without context.
+In the experimental part, we will study whether such a modification brings a significant performance improvement.
 III Janssen spectrogram inpainting (Janssen-TF)
@@
 3.1.2].
-In the inpainting method proposed by Janssen et al.
+In the time-domain inpainting method proposed by Janssen et al.
 [2], the restored audio signal is supposed to follow the AR model (4) while both the missing samples in the time-domain and the unknown AR parameters $\mathbf{a}$ are being estimated.
 Janssen et al.
-approach the problem iteratively by repeating two principal steps in the $i$-th iteration:
+approach the problem iteratively by performing two principal steps in the $i$-th iteration:
   1. 1.
@@
 Regarding step 2, we propose to adapt the Janssen iteration to the TF setting, denoted Janssen-TF.
-In our adaptation, the signal spectrogram is constrained to fit the observation, $\mathbf{X}^{\textup{cor}}$, while the signal follows the estimated AR model.
+In our adaptation, the signal spectrogram is constrained to fit the observation, $\mathbf{X}^{\textup{cor}}$, while the signal itself follows the estimated AR model.
 To formalize this modification, note that the error in (4) can be written as $\mathbf{e}=\mathbf{A}\mathbf{x}$, where $\mathbf{A}\in\mathbb{R}^{(N+p)\times N}$ is a Toeplitz matrix composed using the (fixed) coefficients $\mathbf{a}$.
@@
 $\mathbf{x}^{(i+1)}=\arg\min_{\mathbf{x}}\tfrac{1}{2}\|\mathbf{A}\mathbf{x}\|^{2}+\iota_{\mathit{\Gamma}_{\textup{TF}}}(\mathbf{L}\mathbf{x}),$ (5)
-where $\iota_{\mathit{\Gamma}_{\textup{TF}}}$ denotes the indicator function of the convex set $\mathit{\Gamma}_{\textup{TF}}$ of feasible spectrograms.
+where $\mathbf{L}$ is the STFT and $\iota_{\mathit{\Gamma}_{\textup{TF}}}$ denotes the indicator function of the convex set $\mathit{\Gamma}_{\textup{TF}}$ of feasible spectrograms.
 [^3] Note that the original Janssen algorithm minimizes the function $\frac{1}{2}\|\mathbf{A}\mathbf{x}\|+\iota_{\mathit{\Gamma}_{\textup{T}}}(\mathbf{x})$ in subproblem 2.
@@
 IV Experiments
-IV-A Setup and metrics
-For our experiments, we adopt the audio dataset [^5] https://github.com/fmiotello/dpai/ used for the evaluation of DPAI [18].
+IV-A Data, setup and metrics
+For our first experiment, we adopt the audio dataset [^5] https://github.com/fmiotello/dpai/ used for the evaluation of DPAI [18].
 It consists of 6 music and 2 speech recordings of 5 seconds in length, sampled at 16 kHz.
 Only minor modifications were made to the original distribution of gaps (i.e., the placement of the missing spectrogram columns), such that each test excerpt includes 5 gaps of the same length.
-The gap lengths range from 1 up to 6 missing columns.
-In turn, the entire test set comprises of $8\times 6=48$ corrupted audio signals.
-Even though the dataset is relatively small and includes only a few types of audio signals, we decided to stick to it because its use justifies a fair comparison of the proposed method with DPAI.
-The choice allows the use of exactly the same U-net architecture of DPAI from [18].
-In particular, we choose their “best2” setting.
-The number of epochs of the Adam optimizer is 5000.
-The STFT (i.e., the operator $\mathbf{L}$) utilizes the 2048-sample-long Hann window (which corresponds to 128 ms), with 75% overlap and 2048 frequency channels; this setting is the default for the time-frequency transform of the librosa package [^6] https://librosa.org/doc/0.10.2/generated/librosa.stft.html, as used in [18], and ensures that $\mathbf{L}$ forms a Parseval tight frame.
-To objectively evaluate the reconstruction quality, we assess the difference between the inpainted signal $\hat{\mathbf{x}}$ and a clean reference $\mathbf{x}^{\textup{orig}}$ using two standard metrics:
+The gap lengths range from 1 to 6 missing columns.
+In turn, the entire test set comprises $8\times 6=48$ corrupt spectrograms.
+Even though the dataset is relatively small and includes only a few types of audio signals, it provides a fair comparison of the proposed method with DPAI—the choice allows the use of exactly the same U-net architecture of DPAI from [18].
+In particular, we chose their “best2” setting, and 5000 epochs of the Adam optimizer.
+To validate the results on a larger dataset, we hand-picked 60 music recordings of various complexity from the IRMAS dataset [31, 32].
+To enable the use of DPAI, the recordings were subsampled to 16 kHz and shortened to 5 seconds.
+The STFT (i.e., the operator $\mathbf{L}$) utilizes a 2048-sample-long Hann window (which corresponds to 128 ms), with 75% overlap and 2048 frequency channels; this setting is the default for the time-frequency transform of the librosa package [^6] https://librosa.org/doc/0.10.2/generated/librosa.stft.html, as used in [18], and ensures that $\mathbf{L}$ forms a Parseval tight frame.
+We compare the proposed method Janssen-TF-ADMM with three baseline methods:
+  • DPAI reconstruction with context,
+  • DPAI reconstruction without context,
+  • gapwise Janssen method in the time domain [12].
+The gapwise method is a twist to the original Janssen algorithm [2], which, instead of processing by means of overlapping frames, makes use of a single segment consisting of the whole gap and its nearby context.
+To objectively evaluate the reconstruction quality, we assess the difference between the inpainted signal $\hat{\mathbf{x}}$ and the clean reference $\mathbf{x}^{\textup{orig}}$ using two standard metrics:
   1. 1.
@@
     objective difference grade (ODG) computed using the PEMO-Q [33], which takes psychoacoustics into account and predicts the subjective evaluation of the difference between $\hat{\mathbf{x}}$ and $\mathbf{x}^{\textup{orig}}$ on the scale from $-4$ to 00 (from “very annoying” to “imperceptible”).
-Above that, we run
+In addition, we run
   1. 3.
     a subjective listening test containing a limited number of excerpts.
-The MUSHRA-type test [32] was performed that includes the reference signal, the DPAI reconstruction with context, the DPAI reconstruction without context, the DPAI with context but averaged result from four random initializations, the gapwise Janssen method in the time domain [12], the Janssen-TF-ADMM, and finally the hidden anchor, i.e., the signal reconstructed from the spectrogram with gaps.
-The participants were asked to score the perceived similarity of the signals with the reference on a continuous scale from 0 to 100, with the help of the webMUSHRA environment [^7] https://github.com/audiolabs/webMUSHRA/.
+The MUSHRA-type test [34] was performed, which includes the reference signal, the competing recosntructions, and finally the hidden anchor, i.e., the signal reconstructed from the spectrogram with gaps.
+The participants were asked to score the perceived similarity of the signals with the reference signal on a continuous scale from 0 to 100, with the help of the webMUSHRA environment [^7] https://github.com/audiolabs/webMUSHRA/.
 The test was run in a quiet music studio, using a professional sound card and headphones.
@@
 IV-B Results
-First, we evaluate the proposed modifications of DPAI.
-As seen in Fig.
-1, averaging the realizations brings improvement in quality in terms of SNR, but it comes at the cost of a higher time complexity.
-The differences observed in Fig.
-1 were further tested using the Wilcoxon signed rank test [^8] https://www.mathworks.com/help/stats/signrank.html, which proved that the median of the SNR using DPAI with averaging is significantly higher compared to the original DPAI with context.
-On the other hand, a significant difference in case of the ODG metric has not been proven.
-001010101020202020303030304040404050505050001010101020202020303030304040404050505050SNR (dB), DPAI with contextSNR (dB), DPAI averagedGaplength1 2 3 4 5 6 44-4- 433-3- 322-2- 211-1- 10044-4- 433-3- 322-2- 211-1- 100ODG, DPAI with contextODG, DPAI averaged
-Figure 1: Comparison of DPAI reconstruction with context versus the proposed DPAI variant averaging four random initializations.
-In contrast to the averaging of realizations, the averaging across the final epochs from a single realization does not increase the computational load and brings better quality mostly to audio recordings reconstructed without context—the greatest improvement (averaged over test signals) was 0.16 ODG for the gap of 2 columns and 15 epochs averaged.
-However, this strategy has a negligible impact when considering reconstructions with context.
-The objective comparison that includes the proposed Janssen-TF method is shown in Fig.
-2.
-The first observation is that the time-domain-based Janssen method [12] cannot compete with the other methods.
-This is due to the incompatibility of gaps in the time and time-frequency domains (a sharp mask in the spectrogram domain affects a longer signal portion when transformed into the time domain).
-While the average performance of the DPAI variants mostly lie within the confidence intervals of each other, suggesting limited significance of the differences, the proposed Janssen-TF-ADMM method largely outperforms the competitors in both SNR and ODG (except for the longest gaps).
+The objective comparison, including the proposed Janssen-TF method, is shown in Fig.
+1 for the case of the DPAI dataset.
+The first observation is that the time-domain-based Janssen method [12] is not competitive with the other approaches.
+This can be explained by the incompatibility of gaps in the time and time-frequency domains (a sharp mask in the spectrogram domain affects a longer signal portion when transformed into the time domain, and the whole affected segment is treated as missing).
+While the average performance of the DPAI variants mostly lies within the confidence intervals of each other, suggesting limited significance of the differences, the proposed Janssen-TF-ADMM method largely outperforms the competitors in both SNR and ODG (except the longest gaps).
 Figure 1: Comparison of the inpainting methods using objective metrics – SNR (top) and ODG (bottom).
-The plots show results averaged over the 8 test signals together with the bootstrap interval estimates of the mean values at the $\alpha=5\%$ significance level [33, p.
+The plots show results averaged over the 8 test signals together with the bootstrap interval estimates of the mean values at the $\alpha=5\%$ significance level [35, p.
 160].
 If the intervals do not overlap, it may be concluded that the difference of the means is statistically significant.
-To support the objective evaluation, Fig.
+To support the objective evaluation on the first dataset, Fig.
 2 presents the results of the subjective listening test.
-Its results correlate well with the objective evaluation; in particular, Janssen-TF-ADMM is superior to the competing methods.
+Its results correlate well with the objective evaluation; in particular, Janssen-TF-ADMM is superior to the other methods.
 Moreover, this result is statistically significant in terms of medians.
 On the other hand, the DPAI variants scored almost the same.
-This is in line with the ODG results in Fig.
-1, even though the listening test featured only a selection of the test signals.
-anchorreferenceDPAI without contextDPAI with contextDPAI averagedgapwise JanssenJanssen-TF-ADMM0020202020404040406060606080808080100100100100MUSHRA score
+anchorreferenceDPAI without contextDPAI with contextgapwise JanssenJanssen--TF-ADMM0020202020404040406060606080808080100100100100MUSHRA score
 Figure 2: A boxplot showing the distribution of scores in the listening test.
@@
 The notches (filled areas) around the medians (orange lines) are constructed such that boxes whose notches do not overlap have different medians at the 5% statistical significance level.
-A single run of DPAI takes ca 19 minutes on the NVIDIA Tesla V100S GPU with 32 GB of memory, regardless the number of missing spectrogram columns.
-Computing a single signal with Janssen-TF-ADMM takes 10 to 20 minutes, directly proportional to the number of columns; this corresponds to a PC with Intel Core i7 3.40 GHz processor, 32 GB RAM.
+Finally, Fig.
+3 shows the evaluation on the IRMAS subset comprising 60 signals.
+This evaluation supports the previous analysis and leads to the conclusion that the proposed method is superior to the others.
+For shorter gaps, DPAI with context is preferable over DPAI not exploiting context.
+Figure 3: SNR and ODG results on the IRMAS dataset.
+Regarding computational demands, a single run of DPAI takes ca 19 minutes on the NVIDIA Tesla V100S GPU with 32 GB of memory, regardless of the number of missing spectrogram columns.
+Reconstructing a single signal with Janssen-TF-ADMM takes 10 to 20 minutes, proportional to the number of columns; this performance corresponds to a PC with Intel Core i7 3.40 GHz processor, 32 GB RAM.
 V Conclusion and Outlook
 The paper has shown that the proposed method of spectrogram inpainting, Janssen-TF, performs significantly better than the recently introduced DPAI algorithm, which is based on the deep prior idea.
-This conclusion has been certified both by objective and subjective tests.
-The proposed minor modifications of DPAI led to a slight improvement in terms of the SNR, but have not been reflected in the perceived sound quality.
-In the future, it might be interesting to use the same adaptation strategy with other methods for the time-domain inpainting, such as the PHAIN regularized algorithm [13].
+The conclusion has been certified both by objective and by subjective tests.
 Matlab source codes for Janssen-TF and other supplementary material are publicly available at https://github.com/rajmic/spectrogram-inpainting.

### v3

@@
 Abstract
-The paper focuses on inpainting missing parts of an audio signal spectrogram.
-The autoregression-based Janssen algorithm, the state-of-the-art for the time-domain audio inpainting, is adapted for the time-frequency setting.
-This novel method, termed Janssen-TF, is compared to the deep-prior neural network approach using both objective metrics and a subjective listening test, proving Janssen-TF to be superior in all the considered measures.
+The paper focuses on inpainting missing parts of an audio signal spectrogram, i.e., estimating the lacking time-frequency coefficients.
+The autoregression-based Janssen algorithm, a state-of-the-art for the time-domain audio inpainting, is adapted for the time-frequency setting.
+This novel method, termed Janssen-TF, is compared with the deep-prior neural network approach using both objective metrics and a subjective listening test, proving Janssen-TF to be superior in all the considered measures.
 Index Terms:
-audio inpainting, autoregression, deep prior, time-frequency, spectrogram.
+audio inpainting, autoregression, deep prior, DPAI, time-frequency, spectrogram.
 I Introduction
@@
 If compact regions of such samples are present, they are often referred to as the gaps.
-For the described task, plenty of methods have been developed: those based on the autoregressive character of audio [1, 2], methods based on the sparsity of the time-frequency representation of audio [3, 4, 5, 6], or methods relying on similarities in audio recordings [7, 8, 9].
+For the described task, numerous methods have been developed: those exploiting the autoregressive character of audio [1, 2], methods based on the sparsity of the time-frequency representation of audio [3, 4, 5, 6], or methods relying on similarities in audio recordings [7, 8, 9].
 Quite recently, deep learning audio inpainting methods appeared [10, 11].
@@
 Nonetheless, many practical algorithms, such as classic audio codecs (MP3, AAC, etc.), operate on short-time signal spectra.
-From this perspective, it is valuable to think about inpainting in the time-frequency domain.
-[^1] In other fields, one can go even further and require inpainting purely in the frequency domain, see for example [14], [15].
-In such a case, a portion of time-frequency (TF) coefficients222 typically in the form of a complex-valued matrix are missing, which need to be recovered to allow a later reconstruction of the time-domain signal.
+From this perspective, it is valuable to think about inpainting in the time-frequency domain, [^1] In other fields, one can go even further and require inpainting purely in the frequency domain, see for example [14], [15].
+besides the fact that it is an interesting problem per se.
+In the case under consideration, a portion of time-frequency (TF) coefficients222 typically in the form of a complex-valued matrix is missing, which need to be recovered to allow a later reconstruction of the time-domain signal.
 Indeed, there have been a few methods developed in the TF context, all of them relying on neural modeling [16, 17, 18].
-The most recent work [18] is based on the concept of a “deep prior”, a neural network that does not need training in the most common sense.
-Note that a compact gap in time bears some fundamental differences compared with a gap in the form of a missing number of consecutive columns of a spectrogram.
-This follows from the nature of time-frequency transforms, which use windows with overlaps to compute spectrograms, see for instance [19].
-In particular, a gap in a spectrogram does not fully correspond to a time-domain gap and vice versa; there is always an amplitude fading in one of the domains thanks to the windowing.
+The most recent work [18] is based on the concept of a “deep prior”, a neural network that does not need training data besides the corrupted observation [19].
 The goal of this paper is to adapt the Janssen algorithm [2], proven successful in the time-domain case, for the TF-domain case, and compare its performance with the recent deep-prior-based approach of [18], both using objective metrics and by listening tests.
-II Deep Prior Audio Inpainting (DPAI)
+II Gaps and the Short-time Fourier Transform
+Note that a compact gap in the time domain is far from being equivalent to a gap in the TF domain, which can be associated with the uncertainty principle.
+When a time-domain signal is transformed using the Short-time Fourier Transform (STFT), it is first divided into overlapping frames, which are weighted using windows such as the Hann window (see, e.g., [20]).
+When a gap is present, the number of TF coefficients affected by the gap depends on how much the frames overlap.
+In practical scenarios, the gap overlaps with more than one frame; setting the missing samples to zeros then reduces the value of the corresponding TF coefficients, leading to a kind of TF-fadeout.
+For wide gaps, even a number of completely zero columns can appear in the TF domain (i.e., in the spectrogram), see the analysis in [21].
+Since the inverse STFT also uses windowed frames, the same holds in the opposite direction: consecutive empty spectrogram columns result in a decrease in the time-domain amplitude, governed by the window shape.
+In turn, there is no way how to equitably compare time-domain inpainting techniques with those specialized for the TF-domain.
+III Deep Prior Audio Inpainting (DPAI)
 The deep prior, introduced in [19], is a concept used for the reconstruction of corrupted signals using a convolutional neural network (CNN).
 In contrast to most CNN applications, [19] does not utilize a training dataset; the reconstruction only involves a proper CNN architecture and the corrupted signal.
-The deep prior concept has also proven that the network architecture matters much more than previously thought.
-II-A Application of Deep Prior to Audio Inpainting
+The deep prior concept has also proven that the network architecture matters significantly more than previously thought.
+III-A Application of Deep Prior to Audio Inpainting
 The deep prior reconstruction can be utilized in a variety of applications, as outlined in [19].
@@
 Such a relation naturally characterizes the feasible set $\mathit{\Gamma}_{\textup{T}}=\{\mathbf{x}\,|\,\mathbf{m}\odot\mathbf{x}=\mathbf{m}\odot\mathbf{x}^{\textup{orig}}\}$.
-An encoder–decoder CNN denoted $f_{\theta}(\mathbf{n})$ is assumed to be able to generate any signal of size $N$ using a random but fixed noise input $\mathbf{n}\in\mathbb{R}^{N}$, using a proper $\theta$.
+An encoder–decoder CNN denoted $f_{\theta}(\mathbf{n})$ is assumed to be able to generate any signal of size $N$ using a random but fixed input noise $\mathbf{n}\in\mathbb{R}^{N}$, using a proper $\theta$.
 This is a reliable assumption since the number of network parameters ${\theta}$ (weights and biases) is significantly greater than $N$.
-In the deep prior reconstruction, parameters $\theta$ are initialized randomly. To obtain a reconstruction result, $\theta$ is then optimized as follows:
+In the deep prior reconstruction, also the parameters $\theta$ are initialized randomly. To obtain a reconstruction result, $\theta$ is then optimized as follows:
 $\theta^{*}=\operatorname*{argmin}_{\theta}E(\mathbf{m}\odot f_{\theta}(\mathbf{n}),\mathbf{x}^{\textup{cor}}),$ (1)
-where $E$ is a differentiable data fidelity function measuring the closeness of its two inputs.
+where $E$ is a differentiable data fidelity function measuring the proximity of its two inputs.
 The optimal $\theta^{*}$ can be obtained using an iterative optimizer, e.g., Adam [22].
@@
 Instead, early stopping of the training is desired, resulting in an estimate $\hat{\theta}$ and a corresponding $\hat{\mathbf{x}}=f_{\hat{\theta}}(\mathbf{n})$ [19].
-The authors of [18] used an analog approach to inpaint audio in the transformed domain.
-Through the use of the short-time Fourier transform (STFT), denoted $\mathbf{L}$ in the following, an audio recording $\mathbf{x}$ is considered merely in the TF domain, where it forms a spectrogram matrix $\mathbf{L}\mathbf{x}=\mathbf{X}\in\mathbb{C}^{T\times F}$.
+The authors of [18] used an analog approach to inpaint audio in the time-frequency domain.
+Through the use of the STFT, denoted $\mathbf{L}$ in the following, an audio recording $\mathbf{x}$ is considered merely in the TF domain, where it forms a spectrogram matrix $\mathbf{L}\mathbf{x}=\mathbf{X}\in\mathbb{C}^{T\times F}$.
 Therefore the CNN operates on spectrograms and $\hat{\mathbf{X}}=f_{\hat{\theta}}(\mathbf{N})$ for a random but fixed matrix $\mathbf{N}$.
 Once $\hat{\mathbf{X}}$ has been established, the time-domain solution is obtained by the inverse STFT.
-Regarding the mask, in [18] only the case of completely missing columns of the spectrogram is considered, i.e., for the TF mask $\mathbf{M}\in\{0,1\}^{T\times F}$ it holds that each column is either full of ones or zeros.
+Regarding the mask, in [18] only the case of completely missing columns of the spectrogram is considered, i.e., for the TF mask $\mathbf{M}\in\{0,1\}^{T\times F}$, each of its columns is either full of ones or zeros.
 Nevertheless, a generalization to other TF masks is straightforward.
 The observed spectrogram $\mathbf{X}^{\textup{cor}}$ is finally modeled as $\mathbf{X}^{\textup{cor}}\!=\mathbf{M}\odot\mathbf{X}^{\textup{orig}}$, where $\mathbf{X}^{\textup{orig}}$ corresponds to $\mathbf{x}^{\textup{orig}}$ in the time domain, i.e., $\mathbf{X}^{\textup{orig}}=\mathbf{L}\mathbf{x}^{\textup{orig}}$.
-Note that by analogy with the time-domain case, we have the feasible set $\mathit{\Gamma}_{\textup{TF}}=\{\mathbf{X}\,|\,\mathbf{M}\odot\mathbf{X}=\mathbf{M}\odot\mathbf{X}^{\textup{orig}}\}$.
-The loss function $E$ in (1), which is optimized with respect to $\theta$, is in the form
-$E=\alpha_{1}\mathrm{MSE}(\mathbf{M}\odot\mathbf{X},\mathbf{X}^{\textup{cor}})+\alpha_{2}\mathrm{MSS}(\mathbf{M}\odot\mathbf{X},\mathbf{X}^{\textup{cor}}),$ (2)
-where MSE is the Mean Squared Error and $\mathrm{MSS}$ is the multi-scale spectrogram loss [22]. In [18], the weights $\alpha_{1}$ and $\alpha_{2}$ have values 1 and 0.1 respectively.
-II-B Architecture
-The U-Net is a well-known CNN architecture consisting of several convolutional layers, downsampling (encoder) and upsampling (decoder) operations.
-In certain applications including audio inpainting, the existence of skip connections between the encoder and decoder stages is important.
-The authors of [18] used an architecture similar to U-Net called MultiResUNet [23] with an additional inclusion of the so-called harmonic convolution [24], which proved to be the best among multiple tested architectures in [18].
-The harmonic convolution is a variation on the common 2D convolution that takes into account higher harmonics of audio.
-II-C Variants
-The DPAI reconstructs the whole $\hat{\mathbf{X}}$.
-Yet, considering that only a part of the spectrogram was unavailable, the authors of [18] evaluate the performance of their method after injecting the reliable coefficients back to the solution as a postprocessing step.
+Note that by analogy with the time-domain case, we obtain the feasible set $\mathit{\Gamma}_{\textup{TF}}=\{\mathbf{X}\,|\,\mathbf{M}\odot\mathbf{X}=\mathbf{M}\odot\mathbf{X}^{\textup{orig}}\}$.
+In [18], the loss function $E$ in (1) linearly combines the classic Mean Squared Error between spectrograms and the multi-scale spectrogram loss [23, 24].
+III-B Architecture
+The authors of [18] used an architecture similar to U-Net called MultiResUNet [25].
+It is a common CNN architecture with several layers; the skip connections between the encoder and decoder stages play a significant role in audio inpainting.
+The architecture additionally includes the so-called harmonic convolution [26], which proved to be the best among multiple tested architectures in [18].
+The harmonic convolution is a variation on the common 2D convolution that takes into account higher harmonics in audio.
+III-C Variants
+The DPAI intrinsically reconstructs the whole $\hat{\mathbf{X}}$.
+Yet, considering that only a part of the spectrogram was not observed, the authors of [18] evaluate the performance of their method after injecting the reliable coefficients back to the solution as a postprocessing step.
 Formally, such an action is a projection
@@
 onto the set $\mathit{\Gamma}_{\textup{TF}}$.
-Note that a resulting sharp transition between the regions is smeared in the time domain due to the nature of the STFT [19].
+Note that a resulting sharp transition between the regions is smeared in the time domain due to the nature of the STFT [21]; see also Sec.
+II.
 We call the reconstructions modified by (2) with context, while the others using the entire $\hat{\mathbf{X}}$ are without context.
-In the experimental part, we will study whether such a modification brings a significant performance improvement.
-III Janssen spectrogram inpainting (Janssen-TF)
+In the experimental part, we will study whether employing the context brings an improvement in performance.
+IV Janssen spectrogram inpainting (Janssen-TF)
 A signal $\mathbf{x}=[x_{1},\dots,x_{N}]^{\top}\!\in\mathbb{R}^{N}$ corresponds to an autoregressive process if the following holds:
-$\sum_{i=1}^{p+1}a_{i}x_{n+1-i}=e_{n},\quad n=1,\dots,N+p,$ (4)
-where $\mathbf{a}=[1,a_{2},\dots,a_{p+1}]^{\top}$ is the vector of model parameters (AR coefficients), $p$ is the model order, and $\mathbf{x}$ is zero-padded such that its final length is $N+p$.
-The error term $\mathbf{e}\in\mathbb{R}^{N+p}$ represents a realization of a zero-mean white noise process with variance $\sigma^{2}$ [25, Def.
+$\sum_{i=1}^{p+1}a_{i}x_{n+1-i}=e_{n},\quad n=1,\dots,N+p.$ (3)
+Here, $\mathbf{a}=[1,a_{2},\dots,a_{p+1}]^{\top}$ is the vector of model parameters (AR coefficients), $p$ is the model order, and $\mathbf{x}$ is zero-padded such that its final length is $N+p$.
+The error term $\mathbf{e}\in\mathbb{R}^{N+p}$ represents a realization of a zero-mean white noise process with variance $\sigma^{2}$ [27, Def.
 3.1.2].
@@
     given the model parameters $\mathbf{a}^{(i)}$, estimate a new temporary solution $\mathbf{x}^{(i+1)}$ which fits the observed samples and the norm of the model error $\mathbf{e}$ is minimized.
-Step 1 can be solved efficiently using standard methods, such as the Durbin–Levinson recursion (see e.g.
-the function lpc in Matlab).
-Regarding step 2, we propose to adapt the Janssen iteration to the TF setting, denoted Janssen-TF.
-In our adaptation, the signal spectrogram is constrained to fit the observation, $\mathbf{X}^{\textup{cor}}$, while the signal itself follows the estimated AR model.
+Step 1 can be solved efficiently using standard methods, such as the Durbin–Levinson recursion (see for example the lpc function in Matlab).
+Step 2 boils down to a mean squares problem which can be treated e.g.
+via matrix inversion.
+The proposed adaptation of the Janssen iteration to the TF setting, denoted Janssen-TF, lies in modifying step 2.
+In our case, the signal spectrogram is constrained to fit the observation, $\mathbf{X}^{\textup{cor}}$, at the reliable positions, while the signal itself follows the estimated AR model.
 To formalize this modification, note that the error in (3) can be written as $\mathbf{e}=\mathbf{A}\mathbf{x}$, where $\mathbf{A}\in\mathbb{R}^{(N+p)\times N}$ is a Toeplitz matrix composed using the (fixed) coefficients $\mathbf{a}$.
@@
 where $\mathbf{L}$ is the STFT and $\iota_{\mathit{\Gamma}_{\textup{TF}}}$ denotes the indicator function of the convex set $\mathit{\Gamma}_{\textup{TF}}$ of feasible spectrograms.
-[^3] Note that the original Janssen algorithm minimizes the function $\frac{1}{2}\|\mathbf{A}\mathbf{x}\|+\iota_{\mathit{\Gamma}_{\textup{T}}}(\mathbf{x})$ in subproblem 2.
+[^3] The original Janssen algorithm minimizes the function $\frac{1}{2}\|\mathbf{A}\mathbf{x}\|+\iota_{\mathit{\Gamma}_{\textup{T}}}(\mathbf{x})$ in subproblem 2.
 As such, the problem (4) is convex and it is solvable using standard numerical methods.
-We propose to employ the alternating direction method of multipliers (ADMM), which includes two principal steps [26, Sec.
+We propose to employ the alternating direction method of multipliers (ADMM), which includes two principal steps [28, Sec.
 3.1.1]:
@@
 1.
-For the simplification of (6a), we have posed the assumption [^4] This assumption is met for instance if $\mathbf{L}$ corresponds to a Parseval tight frame, which can be ensured by a suitable setting of the STFT.
-$\mathbf{L}^{*}\mathbf{L}=\mathbf{I}$, leading to $\bar{\mathbf{x}}^{(k+1)}=\operatorname{prox}_{\frac{1}{2}\|\cdot\|^{2}}(\mathbf{L}^{*}(\mathbf{z}^{(k)}-\mathbf{u}^{(k)}))$, see [27, Rem.
+For the simplification of (5a), we have posed the assumption [^4] This assumption is met for instance if $\mathbf{L}$ corresponds to a Parseval tight frame, which can be ensured by a suitable practical setting of the STFT.
+$\mathbf{L}^{*}\mathbf{L}=\mathbf{I}$, leading to $\bar{\mathbf{x}}^{(k+1)}=\operatorname{prox}_{\frac{1}{2}\|\cdot\|^{2}}(\mathbf{L}^{*}(\mathbf{z}^{(k)}-\mathbf{u}^{(k)}))$, see [29, Rem.
 3], which then corresponds to line 1 of Alg.
-1 [28, Table 1].
+1 [30, Table 1].
 Similarly, (5b) boils down to line 1 of the algorithm; this is computed via (2).
@@
 Algorithm 1 Janssen-TF-ADMM
-Note that several algorithms can be used to solve (5) numerically, such as the primal–dual algorithm [29] or the Douglas–Rachford or the forward–backward algorithms [28], combined with the approximal operator of the function $\iota_{\mathit{\Gamma}_{\textup{TF}}}(\mathbf{L}\,\cdot)$ [30].
+Note that several other algorithms can be used to solve (4) numerically, such as the primal–dual algorithm [31] or the Douglas–Rachford or the forward–backward algorithms [30], combined with the approximal operator of the function $\iota_{\mathit{\Gamma}_{\textup{TF}}}(\mathbf{L}\,\cdot)$ [32].
 However, the variant with ADMM described above, denoted Janssen-TF-ADMM, exhibited the best performance out of six algorithms in our preliminary testing.
-IV Experiments
-IV-A Data, setup and metrics
+V Experiments
+V-A Data, setup and metrics
 For our first experiment, we adopt the audio dataset [^5] https://github.com/fmiotello/dpai/ used for the evaluation of DPAI [18].
@@
 To enable the use of DPAI, the recordings were subsampled to 16 kHz and shortened to 5 seconds.
-The STFT (i.e., the operator $\mathbf{L}$) utilizes a 2048-sample-long Hann window (which corresponds to 128 ms), with 75% overlap and 2048 frequency channels; this setting is the default for the time-frequency transform of the librosa package [^6] https://librosa.org/doc/0.10.2/generated/librosa.stft.html, as used in [18], and ensures that $\mathbf{L}$ forms a Parseval tight frame.
+The STFT (i.e., the operator $\mathbf{L}$) utilizes a 2048-sample-long Hann window (corresponding to 128 ms [^6] This exceeds the standard window length for speech processing; however, a change of this hyperparameter would cause the need for changing the DPAI architecture.), with 75% overlap and 2048 frequency channels; this setting is the default for the time-frequency transform of the librosa package [^7] https://librosa.org/doc/0.10.2/generated/librosa.stft.html, as used in [18], and ensures that $\mathbf{L}$ forms a Parseval tight frame.
 We compare the proposed method Janssen-TF-ADMM with three baseline methods:
@@
   • DPAI reconstruction without context,
-  • gapwise Janssen method in the time domain [12].
-The gapwise method is a twist to the original Janssen algorithm [2], which, instead of processing by means of overlapping frames, makes use of a single segment consisting of the whole gap and its nearby context.
+  • gap-wise Janssen method in the time domain [12].
+The gap-wise method is a recent twist to the original Janssen algorithm [2], which is shown in [12] to be the state-of-the-art time-domain inpainter.
+As explained in Sec.
+II, the time-domain Janssen method is not directly comparable with the first two.
+We use it such that we transform the corrupted spectrogram $\mathbf{X}^{\textup{cor}}$ to the time-domain and zero-out all samples affected by the zero columns of $\mathbf{X}^{\textup{cor}}$.
 To objectively evaluate the reconstruction quality, we assess the difference between the inpainted signal $\hat{\mathbf{x}}$ and the clean reference $\mathbf{x}^{\textup{orig}}$ using two standard metrics:
@@
 Following the recommendation [36], 11 assessors out of 18 were selected through post-screening.
-IV-B Results
-The objective comparison, including the proposed Janssen-TF method, is shown in Fig.
-1 for the case of the DPAI dataset.
+V-B Results
+The objective comparison on the DPAI dataset is shown in Fig.
+1.
 The first observation is that the time-domain-based Janssen method [12] is not competitive with the other approaches.
-This can be explained by the incompatibility of gaps in the time and time-frequency domains (a sharp mask in the spectrogram domain affects a longer signal portion when transformed into the time domain, and the whole affected segment is treated as missing).
-While the average performance of the DPAI variants mostly lies within the confidence intervals of each other, suggesting limited significance of the differences, the proposed Janssen-TF-ADMM method largely outperforms the competitors in both SNR and ODG (except the longest gaps).
+While the average performance of the DPAI variants mostly lies within the confidence intervals of each other, suggesting limited significance of the difference, the proposed Janssen-TF-ADMM method largely outperforms the competitors in both SNR and ODG (except the longest gaps).
+11112222333344445555666600101010102020202030303030404040405050505060606060gap lengthSNR (dB)DPAI without contextDPAI with contextgap-wise JanssenJanssen-TF-ADMM
+1111222233334444555566663.53.5-3.5- 3.533-3- 32.52.5-2.5- 2.522-2- 21.51.5-1.5- 1.511-1- 10.50.5-0.5- 0.500gap lengthODGDPAI without contextDPAI with contextgap-wise JanssenJanssen-TF-ADMM
 Figure 1: Comparison of the inpainting methods using objective metrics – SNR (top) and ODG (bottom).
-The plots show results averaged over the 8 test signals together with the bootstrap interval estimates of the mean values at the $\alpha=5\%$ significance level [35, p.
+The plots show results averaged over the 8 test signals together with the bootstrap interval estimates of the mean values at the $\alpha=5\%$ significance level [37, p.
 160].
@@
 Moreover, this result is statistically significant in terms of medians.
-On the other hand, the DPAI variants scored almost the same.
-anchorreferenceDPAI without contextDPAI with contextgapwise JanssenJanssen--TF-ADMM0020202020404040406060606080808080100100100100MUSHRA score
+The DPAI variants scored almost the same.
+anchorreferenceDPAI without contextDPAI with contextgap-wise JanssenJanssen--TF-ADMM0020202020404040406060606080808080100100100100MUSHRA score
 Figure 2: A boxplot showing the distribution of scores in the listening test.
@@
 For shorter gaps, DPAI with context is preferable over DPAI not exploiting context.
+11112222333344445555666600101010102020202030303030404040405050505060606060gap lengthSNR (dB)DPAI without contextDPAI with contextgap-wise JanssenJanssen-TF-ADMM
+1111222233334444555566663.53.5-3.5- 3.533-3- 32.52.5-2.5- 2.522-2- 21.51.5-1.5- 1.511-1- 10.50.5-0.5- 0.500gap lengthODGDPAI without contextDPAI with contextgap-wise JanssenJanssen-TF-ADMM
 Figure 3: SNR and ODG results on the IRMAS dataset.
@@
 Reconstructing a single signal with Janssen-TF-ADMM takes 10 to 20 minutes, proportional to the number of columns; this performance corresponds to a PC with Intel Core i7 3.40 GHz processor, 32 GB RAM.
-V Conclusion and Outlook
+VI Conclusion and Outlook
 The paper has shown that the proposed method of spectrogram inpainting, Janssen-TF, performs significantly better than the recently introduced DPAI algorithm, which is based on the deep prior idea.
 The conclusion has been certified both by objective and by subjective tests.
+In the future, it might be interesting to use the same adaptation strategy with other methods for the time-domain inpainting, such as the PHAIN regularized algorithm [13].
 Matlab source codes for Janssen-TF and other supplementary material are publicly available at https://github.com/rajmic/spectrogram-inpainting.